# A Novel Density Profile for Isothermal Cores of Dark Matter Halos

### v2

@@ I Introduction
Despite the success of the $\Lambda$CDM (cosmological constant $\Lambda$ plus collisionless cold dark matter, CDM) model in explaining the large-scale structure of the universe [e.g. 1, 2], challenges remain in matching simulations to observations of small-scale structures [e.g. 3, 4]. For example, many dwarf galaxies exhibit kpc-size cores in contrast to the universal cuspy density profiles seen in CDM-only simulations (core-cusp problem) [e.g. 5, 6]. The dwarf spheroidal galaxies in the Local Group are found to be systematically less dense compared to the subhalos in CDM-only simulations (the too-big-to-fail problem) [e.g. 7]. Although stellar feedback processes have been shown to alleviate these tensions [e.g. 8, 9], a population of compact dwarf galaxies is missing in simulations of CDM (plus baryons) that can produce DM cores [e.g. 10]. The diversity problem of dwarf galaxy structures is still a challenge for the CDM framework [e.g. 11, 12].
@@
Self-interacting dark matter (SIDM) is an important class of DM models that serves as an alternative to CDM and is well motivated by hidden sector models beyond the standard model [e.g. 13, 14, 15]. SIDM provides promising solutions to many small-scale problems (see [16] for a review). In most cases, DM self-interactions are considered elastic for simplicity and will be the main focus of this paper, although inelastic DM self-interactions have raised some attention recently [e.g. 17, 18, 19, 20, 21]. The elastic collisions between DM particles result in the thermalization of the inner part of the halo and lead to the formation of cores [e.g. 22, 23, 24], which could explain the core-cusp and the too-big-to-fail problems. Meanwhile, the sensitivity of SIDM halo structure to halo concentration and baryon dominance results in more diverse rotation curves compared to the CDM case and is more aligned with observations [e.g. 11, 12, 25]. In recent years, SIDM models with very large self-interaction cross sections at lower velocity ranges have gained popularity. It has implications for the diversity problem of dwarf galaxy structures and the dense compact substructures of galaxy clusters found in strong gravitational lensing observations [e.g. 26]. In this regime, the efficient heat conduction and the negative heat capacity of the SIDM halo result in gravothermal collapse [eg. 27, 28, 29], originally discussed in the context of globular clusters, e.g. [30].
@@
Following the initial thermalization phase and throughout the gravothermal collapse process, until reaching the short-mean-free-path regime when heat conduction is only effective at the surface of the core, the halo core develops characteristic plateaus in both the density and velocity dispersion profiles. The emergence of a central plateau in the density profile, commonly referred to as the “core configuration”, is a key feature of SIDM models and is instrumental in addressing the core-cusp problem, as discussed above. In this context, we also define the “isothermal-core configuration” as a state in which both the density and velocity dispersion remain approximately constant within the central region of the halo. Such behaviors of SIDM have been found in cosmological and isolated N-body simulations [e.g. 31, 22, 23, 24, 32]. In the outer regions of the halo, however, the number density and velocity dispersion of DM particles are significantly smaller than in the core, resulting in negligible scattering interactions over the age of the Universe. As a consequence, the density profile asymptotically approaches the NFW profile [33, 34] commonly observed in CDM.
@@
To characterize the properties of DM halos in simulations, it is common to fit the density profiles with certain functional choices based on the core configuration [e.g. 35, 36, 32, 37, 38, 39]. Efforts to derive analytical density profiles based on the isothermal-core configuration have also been made, either by enforcing the isothermal condition in the Jeans equation [e.g. 40, 41, 25] or within the more complex gravothermal fluid formalism [e.g. 32]. However, despite the advantage of directly capturing the isothermal core and being robust in accommodating baryonic configurations, these profiles lack a closed-form functional expression. Here, we introduce a novel functional form for the density profile that provides a close approximation to the isothermal-core configuration. Section II details the different contemporary density profiles, as well as the motivation and derivation of our profile. We also examine the analytical behavior of each density profile and the corresponding velocity dispersion profiles. Section III inspects the fitting and reconstructing results of the profiles in the context of isolated N-body simulations. We conclude the investigation with a summary in Section IV and discuss potential further applications of the profiles in future SIDM studies.
@@ II.1 Core density profiles
with $\beta\in\left[2;4\right]$ [^1] Similar functional forms are also utilized in [38, 39] among others.. This power-law, often fixed and taking the value of $\beta=4$, results in the density profile that converges almost perfectly to the NFW profile for all radii larger than the core radius. $\rho_{\rm{c}}$ and $r_{\rm{c}}$ represent the characteristic core density and radius, respectively. $r_{\rm{s}}^{\prime}$ is a scale radius, characterizing the NFW tail. This is similar in nature but not to be confused with the initial NFW scale radius.
@@
where $M_{\rm{NFW}}^{\prime}\left(r\right)$ is the NFW mass profile derived from the corresponding NFW scale density $\rho_{\rm{s}}^{\prime}$ and scale radius $r_{\rm{s}}^{\prime}$ [^2] Originally, the Read profile was used to characterize the baryon-induced flat core of DM halos, with an additional exponent $\alpha$ in the $\tanh^{\alpha}{\left(r/r_{\rm{c}}\right)}$ term, controlling the core flatness. However, it has also been used in the context of SIDM as a convenient fitting function [e.g. 25, 32, 38].. The characteristic core density $\rho_{\rm{c}}$ of the Read profile can be recovered as $\rho_{\rm{c}}=1.5\rho_{\rm{s}}^{\prime}r_{\rm{s}}^{\prime}/r_{\rm{c}}$. It must be noted here that the scale density and radius of the Read and other profiles (indicated by the prime notation to be distinguished from the NFW profile parameters) are not necessarily the same as the initial NFW profile parameters, with their values changing throughout the evolution of the halo.
For the Robertson-Fischer and Read profiles, when the condition of $r_{\rm{c}}\ll r_{\rm{s}}^{\prime}$ is satisfied, the density profile asymptotically approach a power-law of $\rho\sim r^{-n}$ in the intermediate region of $r_{\rm{c}}\ll r\ll r_{\rm{s}}^{\prime}$. Here, $n$ is the transition index, taking the value of $n=1$. This index controls the steepness of the transition between the constant-density core ($\rho\sim\text{const}$) and the NFW tail ($\rho\sim r^{-3}$), even in the case of $r_{\rm{c}}\sim r_{\rm{s}}^{\prime}$, where intermediate region becomes ill-defined. It has been demonstrated in [29] that in the gravothermal fluid formalism, a self-similar solution for SIDM halos can be found in which $n=2.19$, provided the collisional mean free path is much larger than the gravitational scale height. This is very similar to the value of $n\simeq 2.5$ observed in simulations [needs citation]. These values, which can also vary throughout the halo evolution, as detailed later in Section III.2, are significantly different from the value of $n=1$ taken by the Read and Robertson-Fischer approaches. Thus, modifications or alternative forms of density profiles are necessary. Accommodating this variation requires the Robertson-Fischer profile to increase the number of degrees of freedom (DOF), taking the form
@@ II.2 Isothermal-core density profile
To construct a density profile that satisfies the isothermal-core configuration, we start from the Jeans equation [e.g. 40, 41, 25]
@@
From Equation 23 and the definition of $I$, we also retrieve the core velocity dispersion of our profile
@@ II.3 Analytical characteristics of density profiles
Figure 2: The velocity dispersion profiles reconstructed from the density profiles presented in Figure 1 using the spherical symmetric Jeans equation (Equation 6). Velocity dispersions and radii are scaled with $\sqrt{G\,\rho_{\rm{c}}\,r_{\rm{c}}^{2}}$ and $r_{\rm{c}}$, respectively. Again, the Read (top left), Robertson-Fischer (top right), Yang (bottom left), and our (bottom right) profiles are shown with three choices of $r_{\rm{s}}/r_{\rm{c}}=3,10,30$ (in pink, green, and blue), as well as three choices of $n=1,1.75,2.5$ (in dotted, dashed, and solid lines) for the relevant profiles. The sub-panels present a zoomed-in view of the velocity dispersion profiles within the expected isothermal region of the core $r\lesssim 3\,r_{\rm{c}}$.
@@ III.1 Simulation data & analysis
Here, $\rho_{i}$ and $\hat{\rho}_{i}(\bm{\theta})$ are the density within radial bin $i$ and its expected value calculated from the analytical profile. The approach here is equivalent to taking the maximum likelihood approach, assuming the SNR of the density measurements remains constant across radial bins. This is typically untrue, as the SNR of the density measurement depends on the number of particles within each radial bin, which can vary by orders of magnitude. However, we want to prioritize the fitting of the halo core and can accept a reasonably good description of the density profile (with the enforced minimum SNR of $20$) rather than a statistically exact solution [^5] Another choice of RSS can be found in [37], where the fitting criterion takes into account the Poisson statistics in isolated N-body simulations. This typically leads to the algorithm prioritizing fitting the outer region over the core, which is not ideal for the investigation of core collapse. The deviations of the results between the two choices of RSS are not significant, except in the values of $r_{\rm{s}}^{\prime}$ or when the halos are deep into core collapse at $T\gtrsim 0.8\,\tau$.. To account for the statistical fluctuation of the density measurements, we employ a similar MC method as detailed above, i.e., perturbing the density profile and analysing the resulting parameter distributions.
@@ III.2 Fitting results
In the current picture, $C=\mathcal{O}\left(1\right)$ often takes the value of $C\simeq 0.75$ to match the results of fluid models [e.g. 17, 32, 45] and N-body simulations [e.g. 46, 17, 19, 47]. Here, we use the value of $C\simeq 0.85$ in order to approximate $\tau$ as the onset of the gravothermal catastrophe, i.e. when the core density and velocity dispersion rapidly diverge while the core radius (and mass) vanishes.
Figure 3 shows the density profile fits and the reconstructed one-dimensional velocity dispersion profiles for three different snapshots of the main halo. The first snapshot, $T\simeq 0.02\,\tau$, corresponds to the earliest phase of core formation. The second, $T\simeq 0.45\,\tau$, represents a typical snapshot during core collapse, where self-similarity is most relevant. The results for this snapshot are representative of the halo’s behavior within the time interval of $T\simeq 0.05\text{--}0.85\,\tau$. The final snapshot, $T\simeq 0.94\,\tau$, shows the halo during the gravothermal catastrophe. The density profiles from each snapshot are fitted using the Read, Robertson–Fischer, Yang, and our profiles. Corresponding velocity dispersion profiles are then reconstructed based on the fitted parameters using the spherically symmetric Jeans equation (Equation 6). For reference, the directly measured core density $\rho_{\rm{core}}$ and velocity dispersion $\sigma_{\rm{core}}$ are also displayed. We find that our profile and the Yang profile generally provide good fits to the halo density structures, while the Read profile deviates significantly after the earliest snapshots. Additionally, although the Robertson–Fischer profile provides a reasonable fit during the earlier evolutionary stages, it fails to accurately capture the density structure’s behavior in the later phases of core collapse. The Yang profile and our profile yield notably similar results in fitting the density structures, with only minor deviations observed in the constant-density cores and the intermediate transition regions, i.e., the regions where density starts to decrease. However, these small deviations result in very different behaviors in the velocity dispersion profiles, most clearly seen in the middle panel of Figure 3, with only our profile closely matching simulations and maintaining consistency with the isothermal-core configuration. Exceptions occur only at the earliest snapshots, where the Yang and other profiles provide a better description of the velocity dispersion structures, and at the later stages of core collapse, where none of the profiles can accurately capture the isothermal core.
@@ IV Conclusions and Discussions
With the exception of the earliest stages of core formation and the late-time deep core collapse regime, our profile demonstrates a high level of agreement with simulation data. As a result of its ability to accurately represent dark matter halos in self-similar scenarios, our profile serves as a promising tool for reducing reliance on computationally expensive SIDM simulations of velocity-independent cross section (VICS) models. The profile can act as a benchmark for comparisons between VICS and more complex interaction models using the empirical formulae detailed in Appendix C. The resulting values of the transition index $n$, derived from the density profile fits, may also serve as a diagnostic for self-similarity in more intricate SIDM models. Additionally, having a functional form also facilitates accurate and efficient calculations of the heat conductivity and luminosity structures of DM halos [e.g. 47]. Finally, initial conditions for simulations entering the deep core collapse regime may be constructed directly from our density profile, following the Eddington sampling procedure described in [43], thus mitigating the need to evolve halos from earlier epochs.
@@ Appendix A Analytical calculation of the core velocity dispersion via the isothermal K-G ratio
$g^{\prime\prime}(\bar{r})$ $=-2n+\frac{4n^{2}}{9}\bar{r}^{2}+\frac{2n}{3}+\frac{2n}{3}+\mathcal{O}(\bar{r}^{2}),$ (18)
@@
Here, the extra term $-\left(3-n\right)^{2}\left(\lambda\bar{r}\right)^{2\gamma}/\bar{r}^{2}$ inside the parentheses has been ignored. It is clear from Equation 22 that as $\bar{r}\rightarrow 0$, unless $\gamma\geq 2$, $\left(\lambda\bar{r}\right)^{\gamma}/\bar{r}^{2}\rightarrow\infty$ and $I_{\rm{c,T25}}\rightarrow 0$. In fact, $\gamma=2$ represents a special case with
@@
$I_{\rm{c,T25}}=\frac{1}{2n+\left(3-n\right)\left(\gamma+1\right)\lambda^{2}\,\bar{r}^{\gamma-2}}+\mathcal{O}\left(\bar{r}^{2}\right).$ (24)

### v3

@@ I Introduction
[…]
@@
Self-interacting dark matter (SIDM) is an important class of DM models that serves as an alternative to CDM and is well motivated by hidden sector models beyond the standard model [e.g. 13, 14, 15]. SIDM provides promising solutions to many small-scale problems (see [16] for a review). In most cases, DM self-interactions are considered elastic for simplicity and will be the main focus of this paper, although inelastic DM self-interactions have raised some attention recently [e.g. 17, 18, 19, 20, 21]. The elastic collisions between DM particles result in the thermalization of the inner part of the halo and lead to the formation of cores [e.g. 22, 23, 24], which could explain the core-cusp and the too-big-to-fail problems. Meanwhile, the sensitivity of SIDM halo structure to halo concentration and baryon dominance results in more diverse rotation curves compared to the CDM case and is more aligned with observations [e.g. 11, 12, 25]. In recent years, SIDM models with very large self-interaction cross sections at lower velocity ranges have gained popularity. It has implications for the diversity problem of dwarf galaxy structures and the dense compact substructures of galaxy clusters found in strong gravitational lensing observations [e.g. 26]. In this regime, the efficient heat conduction and the negative heat capacity of the SIDM halo result in gravothermal collapse [e.g. 27, 28, 29], originally discussed in the context of globular clusters, e.g. [30].
@@
Following the initial thermalization phase and throughout the gravothermal collapse process, until reaching the short-mean-free-path regime when heat conduction is only effective at the surface of the core, the halo core exhibits self-similarity in both the density and velocity dispersion profiles. That is, throughout its evolution, the halo profiles can be characterized by specific values of scale density, radius, and velocity dispersion. This behavior has been well described theoretically [e.g., 29] and confirmed in simulations for both velocity-independent and velocity-dependent self-interacting cross sections [e.g., 31, 32]. Such self-similar nature provides the foundation for adopting particular functional forms to describe the density structure of halos. In this self-similar regime, the halo core develops characteristic plateaus in both the density and velocity dispersion profiles. The emergence of a central plateau in the density profile, commonly referred to as the “core configuration”, is a key feature of SIDM models and is instrumental in addressing the core-cusp problem, as discussed above. In this context, we also define the “isothermal-core configuration” as a state in which both the density and velocity dispersion remain approximately constant within the central region of the halo. Such behaviors of SIDM have been found in cosmological and isolated N-body simulations [e.g. 33, 22, 23, 24, 32]. In the outer regions of the halo, however, the number density and velocity dispersion of DM particles are significantly smaller than in the core, resulting in negligible scattering interactions over the age of the Universe. As a consequence, the density profile asymptotically approaches the NFW profile [34, 35] commonly observed in CDM.
@@
To characterize the properties of DM halos in simulations, it is common to fit the density profiles with certain functional choices based on the core configuration [e.g. 36, 37, 32, 38, 39, 40]. Efforts to derive analytical density profiles based on the isothermal-core configuration have also been made, either by enforcing the isothermal condition in the Jeans equation [e.g. 41, 42, 25] or within the more complex gravothermal fluid formalism [e.g. 32]. However, despite the advantage of directly capturing the isothermal core and being robust in accommodating baryonic configurations, these profiles lack a closed-form functional expression. Here, we introduce a novel functional form for the density profile that provides a close approximation to the isothermal-core configuration. Section II details the different contemporary density profiles, as well as the motivation and derivation of our profile. We also examine the analytical behavior of each density profile and the corresponding velocity dispersion profiles. Section III inspects the fitting and reconstructing results of the profiles in the context of isolated N-body simulations. We conclude the investigation with a summary in Section IV and discuss potential further applications of the profiles in future SIDM studies.
@@ II.1 Core density profiles
with $\beta\in\left[2;4\right]$ [^1] Similar functional forms are also utilized in [39, 40] among others.. This power-law, often fixed and taking the value of $\beta=4$, results in the density profile that converges almost perfectly to the NFW profile for all radii larger than the core radius. $\rho_{\rm{c}}$ and $r_{\rm{c}}$ represent the characteristic core density and radius, respectively. $r_{\rm{s}}^{\prime}$ is a scale radius, characterizing the NFW tail. The prime symbol is used to distinguish this parameter from the initial NFW scale radius $r_{\rm{s}}$. Although $r_{\rm{s}}^{\prime}$ is similar in nature, it should not be confused with $r_{\rm{s}}^{\prime}$.
@@
where $M_{\rm{NFW}}^{\prime}\left(r\right)$ is the NFW mass profile derived from the corresponding NFW scale density $\rho_{\rm{s}}^{\prime}$ and scale radius $r_{\rm{s}}^{\prime}$ [^2] Originally, the Read profile was used to characterize the baryon-induced flat core of DM halos, with an additional exponent $\alpha$ in the $\tanh^{\alpha}{\left(r/r_{\rm{c}}\right)}$ term, controlling the core flatness. However, it has also been used in the context of SIDM as a convenient fitting function [e.g. 25, 32, 39].. The characteristic core density $\rho_{\rm{c}}$ of the Read profile can be recovered as $\rho_{\rm{c}}=1.5\rho_{\rm{s}}^{\prime}r_{\rm{s}}^{\prime}/r_{\rm{c}}$. It must be noted here that the scale density and radius of the Read and other profiles (indicated by the prime notation to be distinguished from the NFW profile parameters) are not necessarily the same as the initial NFW profile parameters, with their values changing throughout the evolution of the halo.
For the Robertson-Fischer and Read profiles, when the condition of $r_{\rm{c}}\ll r_{\rm{s}}^{\prime}$ is satisfied, the density profile asymptotically approach a power-law of $\rho\sim r^{-n}$ in the intermediate region of $r_{\rm{c}}\ll r\ll r_{\rm{s}}^{\prime}$. Here, $n$ is the transition index, taking the value of $n=1$. This index controls the steepness of the transition between the constant-density core ($\rho\sim\text{const}$) and the NFW tail ($\rho\sim r^{-3}$), even in the case of $r_{\rm{c}}\sim r_{\rm{s}}^{\prime}$, where intermediate region becomes ill-defined. It has been demonstrated in [29] that in the gravothermal fluid formalism, a self-similar solution for SIDM halos can be found in which $n=2.19$, provided the collisional mean free path is much larger than the gravitational scale height. This is very similar to the value of $n\simeq 2.5$ observed in simulations (as observed in Section III.2). However, we find that this is a difficult constraint. These values, which can also vary throughout the halo evolution, as detailed later in Section III.2, are significantly different from the value of $n=1$ taken by the Read and Robertson-Fischer approaches. Thus, modifications or alternative forms of density profiles are necessary. Accommodating this variation requires the Robertson-Fischer profile to increase the number of degrees of freedom (DOF), taking the form
@@ II.2 Isothermal-core density profile
To construct a density profile that satisfies the isothermal-core configuration, we start from the Jeans equation [e.g. 41, 42, 25]
@@
From Equation 23 and the definition of II, we also retrieve the core velocity dispersion of our profile
@@ II.3 Analytical characteristics of density profiles
Figure 2: The velocity dispersion profiles reconstructed from the density profiles presented in Figure 1 using the spherically symmetric Jeans equation (Equation 6). Velocity dispersions and radii are scaled with $\sqrt{G\,\rho_{\rm{c}}\,r_{\rm{c}}^{2}}$ and $r_{\rm{c}}$, respectively. Again, the Read (top left), Robertson-Fischer (top right), Yang (bottom left), and our (bottom right) profiles are shown with three choices of $r_{\rm{s}}/r_{\rm{c}}=3,10,30$ (in pink, green, and blue), as well as three choices of $n=1,1.75,2.5$ (in dotted, dashed, and solid lines) for the relevant profiles. The sub-panels present a zoomed-in view of the velocity dispersion profiles within the expected isothermal region of the core $r\lesssim 3\,r_{\rm{c}}$.
@@ III.1 Simulation data & analysis
Here, $\rho_{i}$ and $\hat{\rho}_{i}(\bm{\theta})$ are the density within radial bin $i$ and its expected value calculated from the analytical profile. The approach here is equivalent to taking the maximum likelihood approach, assuming the SNR of the density measurements remains constant across radial bins. This is typically untrue, as the SNR of the density measurement depends on the number of particles within each radial bin, which can vary by orders of magnitude. However, we want to prioritize the fitting of the halo core and can accept a reasonably good description of the density profile (with the enforced minimum SNR of $20$) rather than a statistically exact solution [^5] Another choice of RSS can be found in [38], where the fitting criterion takes into account the Poisson statistics in isolated N-body simulations. This typically leads to the algorithm prioritizing fitting the outer region over the core, which is not ideal for the investigation of core collapse. The deviations of the results between the two choices of RSS are not significant, except in the values of $r_{\rm{s}}^{\prime}$ or when the halos are deep into core collapse at $T\gtrsim 0.8\,\tau$.. To account for the statistical fluctuation of the density measurements, we employ a similar MC method as detailed above, i.e., perturbing the density profile and analyzing the resulting parameter distributions.
@@ III.2 Fitting results
In the current picture, $C=\mathcal{O}\left(1\right)$ often takes the value of $C\simeq 0.75$ to match the results of fluid models [e.g. 17, 32, 46] and N-body simulations [e.g. 47, 17, 19, 48]. Here, we use the value of $C\simeq 0.85$ in order to approximate $\tau$ as the onset of the gravothermal catastrophe, i.e. when the core density and velocity dispersion rapidly diverge while the core radius (and mass) vanishes.
Figure 3 shows the density profile fits and the reconstructed one-dimensional velocity dispersion profiles for three different snapshots of the main halo. The first snapshot, $T\simeq 0.02\,\tau$, corresponds to the earliest phase of core formation. The second, $T\simeq 0.45\,\tau$, represents a typical snapshot during core collapse, where self-similarity is most relevant. The results for this snapshot are representative of the halo’s behavior within the time interval of $T\simeq 0.05\text{--}0.85\,\tau$. The final snapshot, $T\simeq 0.94\,\tau$, shows the halo during the gravothermal catastrophe. The density profiles from each snapshot are fitted using the Read, Robertson–Fischer, Yang, and our profiles. Corresponding velocity dispersion profiles are then reconstructed based on the fitted parameters using the spherically symmetric Jeans equation (Equation 6). For reference, the directly measured core density $\rho_{\rm{core}}$ and velocity dispersion $\sigma_{\rm{core}}$ are also displayed. We find that our profile and the Yang profile generally provide good fits to the halo density structures, while the Read profile deviates significantly after the earliest snapshots. This becomes more apparent in the logarithmic slope presented in Figure 6. Additionally, although the Robertson–Fischer profile provides a reasonable fit during the earlier evolutionary stages, it fails to accurately capture the density structure’s behavior in the later phases of core collapse. The Yang profile and our profile yield notably similar results in fitting the density structures, with only minor deviations observed in the constant-density cores and the intermediate transition regions, i.e., the regions where density starts to decrease. However, these small deviations result in very different behaviors in the velocity dispersion profiles, most clearly seen in the middle panel of Figure 3, with only our profile closely matching simulations and maintaining consistency with the isothermal-core configuration. It should be noted, nonetheless, that even our profile exhibits slight deviations from the simulated velocity dispersion around $r_{\rm{s}}$. The underlying reason can be inferred from the deviations in the logarithmic slope, which also indicate that achieving a more accurate reconstruction of the velocity dispersion profile would require adopting a more complex functional form with additional degrees of freedom. The exceptions for the general goodness of fit observation occur only at the earliest snapshots, where the Yang and other profiles provide a better description of the velocity dispersion structures, and at the later stages of core collapse, where none of the profiles can accurately capture the isothermal core.
@@ IV Conclusions and Discussions
With the exception of the earliest stages of core formation and the late-time deep core collapse regime, our profile demonstrates a high level of agreement with simulation data. As a result of its ability to accurately represent dark matter halos in self-similar scenarios, our profile serves as a promising tool for reducing reliance on computationally expensive SIDM simulations of velocity-independent cross section (VICS) models. The profile can act as a benchmark for comparisons between VICS and more complex interaction models using the empirical formulae detailed in Appendix D. The resulting values of the transition index $n$, derived from the density profile fits, may also serve as a diagnostic for self-similarity in more intricate SIDM models. Additionally, having a functional form also facilitates accurate and efficient calculations of the heat conductivity and luminosity structures of DM halos (an example of which is shown in Appendix E). Finally, initial conditions for simulations entering the deep core collapse regime may be constructed directly from our density profile, following the Eddington sampling procedure described in [44], thus mitigating the need to evolve halos from earlier epochs.
@@ Appendix A Analytical calculation of the core velocity dispersion via the isothermal K-G ratio
$g^{\prime\prime}(\bar{r})$ $=-\frac{2n}{3}+\frac{4n^{2}}{9}\bar{r}^{2}+\mathcal{O}(\bar{r}^{3}),$ (18)
@@
Here, the extra term $\left(3-n\right)^{2}\left(\lambda\bar{r}\right)^{2\gamma}/\bar{r}^{2}$ inside the parentheses has been ignored. It is clear from Equation 22 that as $\bar{r}\rightarrow 0$, unless $\gamma\geq 2$, $\left(\lambda\bar{r}\right)^{\gamma}/\bar{r}^{2}\rightarrow\infty$ and $I_{\rm{c,T25}}\rightarrow 0$. In fact, $\gamma=2$ represents a special case with
@@
$I_{\rm{c,T25}}=\frac{1}{2n+\left(3-n\right)\left(\gamma+1\right)\lambda^{\gamma}\,\bar{r}^{\gamma-2}}+\mathcal{O}\left(\bar{r}^{2}\right).$ (24)